Grade 11General Physics

The position vectors of two bodies of masses 2kg and 3kg are given as 2i+3j+5k and 3i_5j+7k. Find the position vector of centre of mass of the system.

Sharanya
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Xcom => (2*(2) + 2*(3))/(5) => 2 m.
Ycom = (2*3 + 3*5)/5 => 21/5 = > 4.2m
Zcom = (2*5 + 3*7)/5 => 31/5 => 6.2m
Hence the postion vecotr of com of the system is => (2i + 4.2j + 6.2k)